Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability - 3 rolls of a number cube

Expert replies
by Valentino99 » Sun Aug 26, 2012 6:30 am
Hello, its a simple problem but I am not able to spot my fault. Please assist.

Q) what is the probability that, on three rolls of a number cube with faces numbered 1 to 6, at least one of the rolls will be 6?

Solution:
Let A be the event that the number cube rolls a 6.
P(A) = 1/6
P(not A) = 5/6

As it is asked, what is probability of at least one of the rolls to be 6,
we can quickly calculate the probability of no rolls to be 6, which is equal to (5/6)*(5/6)*(5/6) = 125/216.
Hence, probability of at least one of the rolls to be 6 = 1 - 125/216 = 91/216

My doubt:
If I think of calculating the probability in direct way, please tell me where I am making mistake.

We are asked : prob. of at least one of the rolls be a 6 == it means, one roll could be a 6 OR two rolls could be a 6 OR all 3 rolls could be a 6

therefore, ans = P(one roll be a 6) + P(two rolls be a 6) + P(all 3 rolls be a 6)
ans = ((1/6)*(5/6)*(5/6)) + ((1/6)*(1/6)*(5/6)) + ((1/6)*(1/6)*(1/6))
ans = 31/216

Please help me correct my discrepancy between getting the answer as 91/216 and 31/216.
Thanks!
Join the discussion
Source: — Problem Solving |

by Brent@GMATPrepNow » Sun Aug 26, 2012 7:02 am
Valentino99 wrote:Hello, its a simple problem but I am not able to spot my fault. Please assist.

Q) what is the probability that, on three rolls of a number cube with faces numbered 1 to 6, at least one of the rolls will be 6?


We are asked : prob. of at least one of the rolls be a 6 == it means, one roll could be a 6 OR two rolls could be a 6 OR all 3 rolls could be a 6

therefore, ans = P(one roll be a 6) + P(two rolls be a 6) + P(all 3 rolls be a 6)
ans = ((1/6)*(5/6)*(5/6)) + ((1/6)*(1/6)*(5/6)) + ((1/6)*(1/6)*(1/6))
ans = 31/216


Please help me correct my discrepancy between getting the answer as 91/216 and 31/216.
Thanks!
The part in green (above) is perfect. However, the part in blue, is missing some pieces.

Let's focus on: P(one roll is a 6)
You wrote that this probability = (1/6)(5/6)(5/6)
This is not correct. This solution considers only the outcome of {6, not 6, not 6}
You also need to consider the outcomes {not 6, 6, not 6} and {not 6, not 6, 6}

So, P(one roll is a 6) = (1/6)(5/6)(5/6) + (5/6)(1/6)(5/6) + (5/6)(5/6)(1/6)
= 75/216

Similarly, for P(two rolls are 6), your solution of (1/6)(1/6)(5/6) considers only one possible outcome {6, 6, not 6}. You also need to consider {6, not 6, 6} and {not 6, 6, 6}

So, P(two rolls are 6) = (1/6)(1/6)(5/6) + (1/6)(5/6)(1/6) + (5/6)(1/6)(1/6)
= 15/216

Your solution for P(all 3 rolls are 6) is perfect.
P(all 3 rolls are 6) = (1/6)(1/6)(1/6)
= 1/216

So, put everything together to get:
P(at least one 6) = 75/216 + 15/216 + 1/216
= 91/216

I hope that helps.

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by Valentino99 » Sun Aug 26, 2012 7:09 am
:) Yes it does. Thank you!

I gave a calm thought and figured it out too but at the time of the post, I was not able to figure it out..!

Thanks for your reply!!
Join the discussion